What is the slope and y-intercept of the line y = - 2/3x - 12 and how do you know?
Answer by Alan3354(69443)

What is the slope and y-intercept of the line y = - 2/3x - 12 and how do you know?
----------
Slope = -2/3
y-int = -12
------------
y = - 2/3x - 12 is in slope-intercept form, y = mx + b
--> m = -2/3 and b = -12
----------
@ x=0, y = -12 --> b = -12
For each unit increase in x, y changes -2/3 --> m = -2/3